Bobby is 3 years older than Chris. Last year, the sum of their ages was 39. how old is each now?

Answer by multiplier(8)

let x- age of chris at present
(x+3)- age of bobby at present, we add 3 to chris's age bec. he is 3 years younger than bobby
last year the sum of their ages is 39
before we get the sum of their ages which is equal to 39, we must subtract 1 year from their ages
(x-1) - age of chris last year
{(x+3)-1} - age of boby last year
proceeding to addition,
(x-1) + ({x+3}-1)=39
x=19 age of chris at present
x+3=19+3=22 age of bobby at present